OpenStudy (anonymous):

The radius, diameter, or circumference of a circle is given. Find the missing measures. Round to the nearest hundredth if necessary. d = 22.7 mm., r = ? , C = ?

OpenStudy (anonymous):

r=d/2 c=\[2\pi r\] just use the formulas above

OpenStudy (anonymous):

oh so 22.7 divided by 2?

OpenStudy (anonymous):

if 22.7 is the diameter then 22.7/2 will be the radius r

OpenStudy (anonymous):

oh okay thanks :) and how do i find the circumference?